# PlaneRecTR++: Unified Query Learning for Joint 3D Planar Reconstruction and Pose Estimation

### v3

@@ Abstract
3D plane reconstruction from images can usually be divided into several sub-tasks of plane detection, segmentation, parameters regression and possibly depth prediction for per-frame, along with plane correspondence and relative camera pose estimation between frames. Previous works tend to divide and conquer these sub-tasks with distinct network modules, overall formulated by a two-stage paradigm. With an initial camera pose and per-frame plane predictions provided from the first stage, exclusively designed modules, potentially relying on extra plane correspondence labelling, are applied to merge multi-view plane entities and produce 6DoF camera pose. As none of existing works manage to integrate above closely related sub-tasks into a unified framework but treat them separately and sequentially, we suspect it potentially as a main source of performance limitation for existing approaches. Motivated by this finding and the success of query-based learning in enriching reasoning among semantic entities, in this paper, we propose PlaneRecTR++, a Transformer-based architecture, which for the first time unifies all sub-tasks related to multi-view reconstruction and pose estimation with a compact single-stage model, refraining from initial pose estimation and plane correspondence supervision. Extensive quantitative and qualitative experiments demonstrate that our proposed unified learning achieves mutual benefits across sub-tasks, obtaining a new state-of-the-art performance on public ScanNetv1, ScanNetv2, NYUv2-Plane, and MatterPort3D datasets.
@@ 1 Introduction
Immersing in a virtual 3D world involves efficient reasoning of surrounding scenes, whose properties need to be frequently updated. Though tremendous efforts have been paid to create authentic 3D geometry from multi-view image observations or even a single image, a trade-off still exists between the reconstruction quality and efficiency, depending on the underlying scene representations. 3D maps composed of sparse primitives such as point clouds are light-weight to maintain but lack topological structures, while dense geometry like volumetric grids and meshes are computationally intensive to acquire and maintain. In this spectrum, planar representation has been proven to be a reliable alternative, which is compact, efficient, expressive, and generalizable enough to be deployed ubiquitously as well. Therefore, in practice it would always be ideal to infer planar information purely from a video sequence, or even a single image. Being a challenging and yet fundamentally ill-posed computer vision problem, single image plane recovery has been extensively researched and focused so far. Early attempts have been made using image processing techniques to extract low-level primitives such as line segments, vanishing points to extract planar structures from input image [1, 2]. Furthermore, multi-view 3D plane reconstruction [3, 4] are investigated where plane-related constraints are introduced to regularize both camera poses as well as dense geometry, e.g., the well-known Manhattan-world assumption [5, 6].
@@
Figure 1: Previous methods [7, 8, 9] with multiple modules V.S. our single stage model. The 3D visualization of previous pipeline is based on the leading NOPE-SAC [9]. The Green and the Blue frustums show the ground truth and predicted cameras of the first images respectively.The fixed Black frustums show the camera of the second image.
@@
As Convolutional Neural Networks (CNNs) have become the mainstream paradigm to tackle computer vision problems [10, 11] in the past few years, their excellence in performance gradually spreads to plane estimation task. Pioneering works such as PlaneNet [12] and PlaneRCNN [13] propose an efficient solution of piece-wise planar structure recovery from a single image using CNNs in a top-down manner. There have also been bottom-up solutions such as PlaneAE [14] and PlaneTR [15] which obtain plane-level masks by a post-clustering procedure on top of learned deep pixel-wise embeddings. Recently, as another emergent fundamental paradigm, Transformers [16] have made great progress on a wide range of vision tasks [17, 18, 19]. The success of vision Transformers not only comes from its realization of global/long-range interaction across images via attention mechanisms, another important factor is the design of query-based set predictions, initially proposed in Detection Transformer (DETR) [20] to enable reasoning between detected instances and their global context, which has been further proven to be particularly effective in high-level vision tasks such as instance and semantic segmentation [21, 22], video panoptic segmentation [23, 24], etc. PlaneTR [15] is one early attempt of using such ideas in single-view image plane reconstruction. It is also inspired by structured-guided learning to integrate additional geometric cues like line segments during training, leading to state-of-the-art performance on the ScanNetv1 dataset and the unseen NYUv2-Plane dataset. Based on high-quality monocular plane predictions, following works have explored their extension to establish a plane-aware structure from motion framework, involving across-view plane correspondence learning and camera pose estimation [7, 25, 8, 9].
@@
However, all of the above mentioned single-/multiview plane recovery methods somewhat dis-entangle the prediction of principle components required for plane reconstruction. As to single-view approaches, PlaneRCNN [13] learns to predict plane offsets from a monocular depth prediction branch while other attributes such as plane mask and normal are estimated separately from colour images, PlaneTR [15] also predicts a monocular depth map apart from the Transformer module, which is later used for acquiring plane segmentation masks using clustering associate embedding [14]. For multi-view ones, SparsePlanes [7], PlaneFormers [8] and NOPE-SAC [9] inherit the above multi-step method as the monocular plane predictor and further adopt a two-stage framework. Initially, they compute planes of 2 frames in the same coordinate system (from the predicted initial pose). Subsequently, complex manual design optimization algorithms [7] or additional neural modules [8, 9] are employed to achieve plane correspondence and refine pose hypotheses by incorporating externally provided camera states and explicit correspondence ground truth. Accordingly, these closely related prediction tasks are usually interleaved, and none of them manage to unify multi-view plane recovery within a single compact model. We conjecture this could be one performance bottleneck for existing data-driven approaches.
@@
Motivated by this finding, we seek to borrow recent advance in query-based learning and aim to design a single, compact and unified model to jointly learn all plane-related tasks, and we expect such design would achieve a mutual benefits among tasks and advance the existing performance of both monocular and multi-view plane segmentation and reconstruction. Extensive experiments results on the ScanNet, the MatterPort3D and the NYUv2-Plane benchmark datasets shows that without using any external priors [13, 15, 7, 8, 9] during training, our unified querying learning model, named PlaneRecTR++, achieves new state-of-the-art performance with a concise structure. Additionally, we have also found that such framework can implicitly discover the spatial plane correspondences to enable precise 3D plane reconstruction.
@@
  • We proposed a first single unified framework to address the challenging 3D plane recovery task, where all closely related sub-tasks are jointly optimized and inferred in a multi-task manner motivated by query-based learning.
  • Without any external priors and supervisions other than input images, we proposed a novel plane-aware attention structure to first tackle sparse view plane reconstruction in a purely end-to-end manner, producing robust cross-view plane correspondences and camera poses.
  • Our proposed method achieves significant gains in terms of model performance and model compactness. Extensive numerical and visual comparisons on four public benchmark datasets demonstrate state-of-the-art performance of our proposed unified query learning, taking full advantages of plane-related cues to achieve mutual benefits.
A previous version of this work was published at ICCV2023 [26]. This paper extends the conference version with the following new contributions. First, to amplify the effectiveness of unified query learning in plane reconstruction, we extend PlaneRecTR [26] to jointly tackle multi-view plane recovery and camera pose estimation in a purely end-to-end manner, i.e. PlaneRecTR++, enabling superior performance on the ScanNetv2 and MatterPort3D datasets. Second, we propose a plane-aware cross attention module to implicitly learn plane correspondences, achieving mutual benefits without requiring pose initialization and correspondence labelling. Third, we conduct extensive and comprehensive experiments with detailed ablation analysis to provide a thorough understanding of PlaneRecTR++, using four public benchmark datasets on both single-view and multi-view plane recovery tasks.
@@ 2 Related Work
Recovering monocular planes empowers 3D planar reconstruction and structural scene understanding. Traditional methods for 3D plane recovery from a single image often rely on strong assumptions of scenes [1, 27] (e.g., the Manhattan world assumption), or require manual extraction of primitives [1, 27, 2], such as superpixels, points and line segments, which may not be applicable to complex actual scenes.
@@
Therefore, existing advanced methods, whether based on direct CNN prediction or embedding clustering, still divide the whole 3D plane recovery task into several steps. In contrast, our plane query learning-based approach offers a unified solution for the aforementioned sub-tasks within the intra-frame component and can be seamlessly extended to address inter-frame requirements in an end-to-end manner.
@@
SparsePlanes [7] is the first learning-based approach for planar reconstruction and relative camera pose estimation from sparse views. Specifically, with monocular planes and their embeddings from PlaneRCNN [13], a dense pixel attention network is applied to estimate a number of $K$=1024 pose hypotheses, all of which are used for plane matching and a final pose via a complex two-step optimization. Furthermore, PlaneFormers [8] utilizes predicted monocular planes and top $h$=9 pose hypotheses from SparsePlanes [7] as input, but replace the intricate handcrafted optimization by $h$ leanable PlaneFormer modules, mitigating the complex optimization issue faced by [7]. NOPE-SAC [9] improves monocular plane quality by replacing PlaneRCNN [13] with a modified PlaneTR [15], and achieves an initial coarse pose from direct regression using the similary pixel attention [7]. It also introduces differentiable optimal transport [31] for plane matching and proposes one-plane pose hypotheses based on corresponding plane pairs, effectively resolving conflicts between numerous pose hypotheses and limited 3D plane correspondences during SparsePlanes’ pose refinement.
However, both PlaneFormers and the leading Nope-SAC still adopt a multi-stage pipeline derived from SparsePlanes, requiring bootstrapping from external priors like initial pose estimation and correspondence supervision. The reason behind this lies in the existence of a chicken-and-egg relationship between explicitly learning camera pose and plane correspondence. In all previous methods [7, 8, 9], the capability of plane embeddings learned by ground truth correspondence supervision is insufficient to accurately track the same plane instance under sparse views. To address this issue, an additional neural model is necessary to provide the initial pose(s), so that monocular planar information could be merged under a unified coordinate system, thereby assisting the original plane embedding in enhancing matching accuracy and ultimately refining the initial pose. Our proposed PlaneRecTR++ deviates from such dilemma using unified plane query learning, actively inferring plane estimations, correspondences, and camera pose in a single-stage framework, without relying on initial pose guidance and supervision for plane matching.
@@ 2.3 Correspondence and Camera Pose Estimation
Camera pose estimation between adjacent images is a fundamental step in multi-view 3D reconstruction[32]. Early studies focused primarily on extracting sparse keypoint correspondences [33, 34] to compute the essential matrix using the five-point solver [35]. Subsequently, significant efforts have been devoted to learning-based approaches for robust keypoint detection [36, 37, 38] and matching [31, 39, 40]. However, pose computed solely from the essential matrix lacks a real scaling factor and causes further challenges in the sparse view setting, where only a limited number of correct correspondences could be found.
@@
Our approach draws inspiration from attention-based modules [31, 39, 25], but further empowers the standard cross attention by a plane-aware attentive design. With our learned plane embeddings as the only input for pose prediction, our method implicitly learns genuine plane correspondences and is able to recover precise camera poses within a compact module.
Figure 2: Overview of our proposed PlaneRecTR++.
@@ 3 PlaneRecTR++ Overview
Our PlaneRecTR++ is an end-to-end unified query learning architecture, designed for the challenging task of joint planar reconstruction and relative camera pose estimation. In Figure 2, the only input to PlaneRecTR++ are two RGB images $I_{1}$ and $I_{2}$ of different views, with their learnable queries serving as the bridge jointly connecting distinct sub-tasks. Specifically, plane queries guide the learning of unified plane embeddings $\mathcal{E}_{\text{plane}}$ for each possible plane candidate within input frames, whose mutual interactions enable their decoding to final planar attributes, cross-view plane correspondences, and relative camera pose. Please note that the above results are obtained by PlaneRecTR++ in one shot without any external priors like initial poses or planes [7, 8, 9].
To better dissect our method, we partition PlaneRecTR++ into two components based on the learning scope: intra-frame and inter-frame plane query learning. The intra-frame plane query learning component aims to recover per-view 3D planes. It unifies various sub-tasks of monocular plane recovery including plane detection, parameter prediction, segmentation and depth estimation, while eliminating the need for multi-step prediction in previous methods [7, 8, 9]. Such design philosophy brings not only simplicity and compactness of overall framework but also mutual benefits among closely related tasks. Furthermore, in the inter-frame plane query learning component, we integrate these learned unified plane embeddings from different views to form an attentive planar association matrix to approximate plane correspondences and capture fetures of paired planes for directly pose regression. This design choice implicitly motivates multi-view consistency of planar embeddings and is proven to be sufficient for precise plane instance tracking, without any direct supervision. As a result, PlaneRecTR++ manages to unify separate multi-stage tasks required in the previous sparse views pipelines [7, 8, 9], including monocular plane recovery, pose initialization, plane matching, and pose refinement. This joint optimization of all sub-tasks naturally enhances efficiency and final performance.
The training process of PlaneRecTR++ also comprises two phases: 1) a monocular pre-training phase, where monocular RGB images are utilized to train the intra-frame component for single-view plane recovery; 2) a joint training phase, where paired RGB images are used to train the complete pipeline of PlaneRecTR++, optimizing both components in an end-to-end manner for planar reconstruction and camera pose estimation on challenging sparse-view datasets. Please note that our framework could also be trained and converged well from scratch (i.e., only applying the joint training phase), however, we found the two-phase training achieves better overall performance without losing its virtue in end-to-end unified query learning. Therefore during experiments we stick to the two-phase training unless otherwise mentioned.
@@
Figure 3: Overview of intra-frame plane query learning (PlaneRecTR). Our model consists of three main modules: (1) A pixel-level module to extract dense pixel-wise image features; (2) A Transformer module to jointly predict 4 plane-related properties from each plane query, including plane classification probability, plane parameter, mask and depth embedding; (3) A plane-level module to calculate dense plane-level binary masks/depths, then filter non-plane predictions and produce the final 3D plane recovery.

### v4

@@ Abstract
The challenging task of 3D planar reconstruction from images involves several sub-tasks including frame-wise plane detection, segmentation, parameter regression and possibly depth prediction, along with cross-frame plane correspondence and relative camera pose estimation. Previous works adopt a divide and conquer strategy, addressing above sub-tasks with distinct network modules in a two-stage paradigm. Specifically, given an initial camera pose and per-frame plane predictions from the first stage, further exclusively designed modules relying on external plane correspondence labeling are applied to merge multi-view plane entities and produce refined camera pose. Notably, existing work fails to integrate these closely related sub-tasks into a unified framework, and instead addresses them separately and sequentially, which we identify as a primary source of performance limitations. Motivated by this finding and the success of query-based learning in enriching reasoning among semantic entities, in this paper, we propose PlaneRecTR++, a Transformer-based architecture, which for the first time unifies all tasks of multi-view planar reconstruction and pose estimation within a compact single-stage framework, eliminating the need for the initial pose estimation and supervision of plane correspondence. Extensive quantitative and qualitative experiments demonstrate that our proposed unified learning achieves mutual benefits across sub-tasks, achieving a new state-of-the-art performance on the public ScanNetv1, ScanNetv2, NYUv2-Plane, and MatterPort3D datasets. Codes are available at https://github.com/SJingjia/PlaneRecTR-PP.
@@ 1 Introduction
Immersing in a virtual 3D world involves efficient reasoning of surrounding scenes, whose properties need to be frequently updated. Though tremendous efforts have been devoted to create authentic 3D geometry from multi-view image observations or even a single image, a trade-off still exists between reconstruction quality and efficiency, depending on the underlying scene representations. 3D maps composed of sparse primitives such as point clouds are light-weight to maintain but lack topological structures, while dense geometry like volumetric grids and meshes are computationally intensive to acquire and maintain. In this spectrum, planar representation has been proven to be a reliable alternative, which is compact, efficient, expressive, and generalizable enough to be deployed ubiquitously as well. Therefore, in practice it would always be ideal to infer planar information purely from a video sequence, or even a single image. Being a challenging and yet fundamentally ill-posed computer vision problem, single image plane recovery has been extensively researched and focused so far. Early attempts have been made using image processing techniques to extract low-level primitives such as line segments, vanishing points to extract planar structures from an input image [1, 2]. Furthermore, multi-view 3D plane reconstruction [3, 4] are investigated where plane-related constraints are introduced to regularize both camera poses as well as dense geometry, e.g., the well-known Manhattan-world assumption [5, 6].
[…]
@@
Figure 1: Previous methods [7, 8, 9] with multiple modules V.S. our single stage model. The 3D visualization of previous pipeline is based on the leading NOPE-SAC [9]. The Green and the Blue frustums show the ground truth and predicted cameras of the first image respectively.The fixed Black frustums show the camera of the second image.
@@
As Convolutional Neural Networks (CNNs) have become the mainstream paradigm to tackle computer vision problems in the past few years [10, 11], their excellence in performance gradually spreads to plane estimation task. Pioneering works such as PlaneNet [12] and PlaneRCNN [13] propose an efficient solution of piece-wise planar structure recovery from a single image using CNNs in a top-down manner. There have also been bottom-up solutions such as PlaneAE [14] and PlaneTR [15] which obtain plane-level masks by a post-clustering procedure on top of learned deep pixel-wise embeddings. Recently, as another emergent fundamental paradigm, Transformers [16] have made great progress on a wide range of vision tasks [17, 18, 19]. The success of vision Transformers not only comes from its realization of global/long-range interaction across images via attention mechanisms, another important factor is the design of query-based set predictions, initially proposed in Detection Transformer (DETR) [20] to enable reasoning between detected instances and their global context, which has been further proven to be particularly effective in high-level vision tasks such as instance and semantic segmentation [21, 22], video panoptic segmentation [23, 24], etc. PlaneTR [15] is one early attempt of using such ideas in single-view image plane reconstruction. It is also inspired by structure-guided learning to integrate additional geometric cues like line segments during training, leading to state-of-the-art performance on the ScanNetv1 dataset and the unseen NYUv2-Plane dataset. Based on high-quality monocular plane predictions, following works have explored their extension to establish a plane-aware structure from motion framework, involving across-view plane correspondence learning and camera pose estimation [7, 25, 8, 9].
@@
However, all of the above mentioned single/multi-view plane recovery methods somewhat disentangle the prediction of principle components required for plane reconstruction. For single-view methods, PlaneRCNN [13] learns to predict plane offset from a monocular depth prediction branch while other attributes such as plane mask and normal are estimated separately from colour images, PlaneTR [15] also predicts a monocular depth map apart from the Transformer module, which is later used for acquiring plane segmentation masks using clustering associate embedding [14]. For multi-view ones, SparsePlanes [7], PlaneFormers [8] and NOPE-SAC [9] inherit the above multi-step methods as the monocular plane predictor and further adopt a two-stage framework. Initially, they compute planes of two frames in the same coordinate system (from the predicted initial pose). Subsequently, complex manual design optimization algorithms [7] or additional neural modules [8, 9] are employed to achieve plane correspondence and refine pose hypotheses by incorporating externally provided camera states and explicit correspondence ground truth. Accordingly, these closely related prediction tasks are usually interleaved, and none of them successfully unifies multi-view plane recovery within a single compact model. We conjecture this could be one performance bottleneck for existing data-driven approaches.
@@
Motivated by this finding, we seek to borrow recent advance in query learning and aim to design a single, compact and unified model to jointly learn all plane-related tasks, and we expect such design would achieve a mutual benefits among tasks and advance the existing performance of both monocular and multi-view plane segmentation and reconstruction. Extensive experiments results on the ScanNet, MatterPort3D and NYUv2-Plane benchmark datasets show that without using any external priors [13, 15, 7, 8, 9] during training, our unified querying learning model, named PlaneRecTR++, achieves new state-of-the-art performance with a concise structure. Additionally, we have also found that such framework can implicitly discover the spatial plane correspondences to enable precise 3D plane reconstruction.
@@
  • We propose a first single unified framework to address the challenging 3D plane recovery task, where all closely related sub-tasks are jointly optimized and inferred in a multi-task manner motivated by query-based learning.
  • Without any external priors and supervisions other than input images, we propose a novel plane-aware attention structure to first tackle sparse view plane reconstruction in a purely end-to-end manner, producing robust cross-view plane correspondences and camera poses.
  • Our proposed method achieves significant gains in terms of model performance and compactness. Extensive numerical and visual comparisons on four public benchmark datasets demonstrate state-of-the-art performance of our proposed unified query learning, taking full advantages of plane-related cues to achieve mutual benefits.
A previous version of this work was published at ICCV2023 [26]. This paper extends the conference version with the following new contributions. First, to amplify the effectiveness of unified query learning in plane reconstruction, we extend PlaneRecTR [26] to jointly tackle multi-view plane recovery and camera pose estimation in a purely end-to-end manner, i.e. PlaneRecTR++, enabling superior performance on the ScanNetv2 and MatterPort3D datasets. Second, we propose a plane-aware cross attention module to implicitly learn plane correspondences, achieving mutual benefits without requiring pose initialization and correspondence labelling. Third, we conduct extensive and comprehensive experiments with detailed ablation analysis to provide a thorough understanding of PlaneRecTR++. The plane embeddings guided by query learning from PlaneRecTR++ not only retain comparable single-view plane recovery capability to [26], but also exhibit superior cross-view consistency, enabling precise plane matching and pose inference.
@@ 2 Related Work
Recovering monocular planes enables 3D planar reconstruction and structural scene understanding. Traditional methods often rely on strong assumptions of scenes [1, 27] (e.g., the Manhattan world assumption), or require manual extraction of primitives [1, 27, 2], such as superpixels and line segments, which may not be applicable to complex real-world scenes.
@@
Therefore, existing advanced methods, whether based on direct CNN prediction or embedding clustering, still divide the whole 3D plane recovery task into several steps. In contrast, our plane query learning offers a unified solution for the aforementioned sub-tasks within the intra-frame component and can be seamlessly extended to address inter-frame requirements in an end-to-end manner.
@@
SparsePlanes [7] is the first learning-based approach for planar reconstruction and pose estimation from sparse views. With monocular planes derived from PlaneRCNN [13] and $1024$ pose hypotheses estimated from a dense pixel attention network, SparsePlanes employs them for a complex two-step optimization to calculate plane correspondences and a final pose. Furthermore, PlaneFormers [8] utilizes predicted monocular planes and top $9$ pose hypotheses from SparsePlanes as input, and replaces the handcrafted optimization by $9$ learnable planeformer modules, mitigating the intricate optimization issue [7]. NOPE-SAC [9] improves monocular plane quality by replacing PlaneRCNN [13] with a modified PlaneTR [15], and achieves an initial coarse pose from direct regression using the similar pixel attention [7]. It also introduces differentiable optimal transport [31] for plane matching and proposes one-plane pose hypotheses based on corresponding plane pairs, resolving conflicts between numerous pose hypotheses and limited 3D plane correspondences during SparsePlanes’ pose refinement.
However, both PlaneFormers and the leading NOPE-SAC still adopt a multi-stage pipeline derived from SparsePlanes, requiring bootstrapping from external initial pose and correspondence supervision. The reason behind this lies in the existence of a chicken-and-egg relationship between explicitly learning camera pose and plane correspondence. In all previous methods [7, 8, 9], the capability of plane embeddings learned by ground truth correspondence supervision is still insufficient to accurately track the same plane instance under sparse views. An additional model is necessary to provide the initial pose(s), so that monocular planes could be merged under a unified coordinate system, thereby assisting the original plane embedding in enhancing matching accuracy and ultimately refining the initial pose. Our proposed PlaneRecTR++ deviates from such dilemma using unified plane query learning, actively inferring plane estimations, correspondences, and camera pose in a single-stage framework, without relying on initial pose guidance and supervision for plane matching.
@@ 2.3 Correspondence and Camera Pose Estimation
Camera pose estimation between adjacent images is a fundamental step in multi-view 3D reconstruction[32]. Early studies focused primarily on extracting sparse keypoint correspondences [33, 34] to compute the essential matrix using the five-point solver [35]. Subsequently, significant efforts have been devoted to learning-based approaches for robust keypoint detection [36, 37, 38] and matching [31, 39, 40]. However, pose computed solely from the essential matrix lacks a real scale and causes further challenges in the sparse view setting, where only a limited number of correct correspondences could be found.
@@
Our approach draws inspiration from attention-based modules [31, 39, 25], but further extends the standard cross attention by a plane-aware attentive design. With our learned plane embeddings as the only input for pose prediction, our method implicitly learns genuine plane correspondences and is able to recover precise camera poses within a compact module.
Figure 2: Overview of the proposed PlaneRecTR++. Our end-to-end model couples intra-frame and inter-frame components through learnable plane queries, unifying numerous interdependent and mutually constraining sub-tasks. Within a single forward pass from input images, PlaneRecTR++ accomplishes joint plane reconstruction and camera pose estimation without initial pose prior.
@@ 3 PlaneRecTR++ Overview
Our PlaneRecTR++ is an end-to-end unified query learning architecture, designed for the challenging task of joint planar reconstruction and relative camera pose estimation. In Figure 2, the only input to PlaneRecTR++ are two RGB images $I_{1}$ and $I_{2}$ of different views, with their learnable queries serving as the bridge jointly connecting distinct sub-tasks. Specifically, plane queries guide the learning of unified plane embeddings $\mathcal{E}_{\text{plane}}$ for each possible plane candidate within input frames, whose mutual interactions enable their decoding to final planar attributes, cross-view plane correspondences, and relative camera pose. Please note that the above results are obtained by PlaneRecTR++ in one shot without external pose prior or correspondences supervision [7, 8, 9].
To better dissect our method, we partition PlaneRecTR++ into two components based on the learning scope: intra-frame and inter-frame plane query learning. The intra-frame component aims to recover per-view 3D planes. It unifies various sub-tasks of monocular plane recovery including plane detection, parameter prediction, segmentation and depth estimation, while eliminating the need for multi-step prediction in previous methods [7, 8, 9]. Such design philosophy brings not only simplicity and compactness of overall framework but also mutual benefits among closely related tasks. Furthermore, in the inter-frame component, we integrate these learned unified plane embeddings from different views to form an attentive planar association matrix to approximate plane correspondences and capture features of paired planes for directly pose regression. This design implicitly motivates multi-view consistency of plane embeddings and is proven to be sufficient for precise plane tracking, without any direct supervision. As a result, PlaneRecTR++ manages to unify separate multi-stage tasks required in the previous sparse views pipelines [7, 8, 9], including monocular plane recovery, pose initialization, plane matching, and pose refinement. This joint optimization of all sub-tasks naturally enhances efficiency and final performance.
The training process of PlaneRecTR++ also comprises two phases: (1) A monocular pre-training phase, where monocular images are utilized to train the intra-frame component for single-view plane recovery; (2) A joint training phase, where paired images are used to train the complete PlaneRecTR++, optimizing both components in an end-to-end manner for planar reconstruction and pose estimation on challenging sparse-view datasets. Please note that our method could also be trained and converged well from scratch, i.e., only applying phase (2), however, we find the two-phase training achieves better overall performance without losing its virtue in end-to-end unified query learning. Therefore we stick to the two-phase training unless otherwise mentioned.
@@
Figure 3: Overview of intra-frame plane query learning (PlaneRecTR). Our intra-frame component consists of three main modules: (1) A pixel-level module to extract dense pixel-wise image features; (2) A Transformer module to jointly predict 4 plane-related properties from each plane query, including plane classification probability, plane parameter, mask and depth embedding; (3) A plane-level module to calculate dense plane-level binary masks/depths, then filter non-plane predictions and produce the final 3D plane recovery.